OpenStudy (anonymous):

If 12 cows produce 70 gallons of milk, how many gallons of milk would 42 cows produce?

OpenStudy (anonymous):

cross multiply

OpenStudy (anonymous):

sory thats wrong

OpenStudy (texaschic101):

12/70 = 42/x (12 cows to 70 gallons = 42 cows to x gallons) cross multiply (12)(x) = (42)(70) 12x = 2940 x = 2940/12 x = 245

OpenStudy (anonymous):